1/1

William Butler Yeats,

He Wishes for the Cloths of Heaven

Had I the heavens’ embroidered cloths, Enwrought with golden and silver light, The blue and the dim and the dark cloths Of night and light and the half light, I would spread the cloths under your feet: But I, being poor, have only my dreams; I have spread my dreams under your feet; Tread softly because you tread on my dreams.

Author of this English poem is William Butler Yeats (∗1865-06-13 – †1939-01-28).

Arg-1104-2229 (2014-12-19 18:39:21)

Pri la poeto vidu la vikipediejon http: // eo. wikipedia. org/ wiki/ William_Butler_Yeats .

http://www.poezio.net/version?poem-id=1104&version-id=,2229